OpenStudy (anonymous):

How to find the zeroes of a function? y = x^3 – 7x^2 – 18x

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

or @ganeshie8 Could either of you help me?

OpenStudy (anonymous):

ok I just figured this out a lil bit so lets try.

OpenStudy (solomonzelman):

set it equal to zero, and factor out of x.

OpenStudy (anonymous):

y= f(x) = x^3 – 7x^2 – 18x = 0 x(x^2 – 7x – 18) = 0 x = 0 (first root) x^2 - 7x - 18 = 0 put this one in quadratic formula to get other roots :D

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf y = x^3 – 7x^2 – 18x }\) \(\LARGE\color{blue}{ \bf x^3 – 7x^2 – 18x =0 }\) \(\LARGE\color{blue}{ \bf x(x^2 – 7x – 18) =0 ~~~~~~~~~~~~x=0}\) \(\LARGE\color{blue}{ \bf x^2 – 7x – 18 =0 }\) \(\LARGE\color{blue}{ \bf (x+2)(x-9) =0 ~~~~~~~x=2,-9}\)

OpenStudy (anonymous):

Kitty solved you the first part let me know if you have problem solving the other part

OpenStudy (solomonzelman):

I mean -2 and 9

OpenStudy (solomonzelman):

So, 0, -2, 9

OpenStudy (anonymous):

Thanks both of you! I could have gotten the rest myself @SolomonZelman but thank you!

OpenStudy (anonymous):

Welcome :D I know right I'm a cute kitty :3 *roof roof* lol

OpenStudy (solomonzelman):

RoseDryer , not a problem!

OpenStudy (anonymous):

So for this one It would be y = x(x + 6) x(x+6)=0 x=0 x+6=0 x=-6 x=0,-6?

OpenStudy (solomonzelman):

Yes, you are indeed correct.

OpenStudy (anonymous):

Great!

OpenStudy (solomonzelman):

Yes :)

OpenStudy (anonymous):

y = (x + 1)(x – 1)(x – 2) Do I just set each one equal to 0?

OpenStudy (solomonzelman):

YES basically, b/c if (x + 1)(x – 1)(x – 2)=0 then one of these (x+1) , (x-1) or (x-2) is zero.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

For this one do I divide by 3 y = 3x3 – 3x and get y=x^3-x?

OpenStudy (anonymous):

try 3x (x^2-1)=0 3x=0 , x^2-1=0 x=0 , x^2 + 0x - 1=0 apply quadratic formula and get other two roots

OpenStudy (solomonzelman):

\[y = 3x^{3} – 3x \]\[ 3x^{3} – 3x =0\]\[ 3(x^{3} –x )=0\]\[x^{3} –x =0\]\[x^{3} =x\]\[x=~~±~1,0\]